OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4

OpenStudy (abhisar):

Hey there! Welcome to OpenStudy :)

OpenStudy (michele_laino):

please, perform the multiplications indicated, namely calculate: -2(x+3)=...

OpenStudy (anonymous):

Hi!

OpenStudy (michele_laino):

@CoDsTeR9000 please, you have to apply the distributive property of multiplication over addition

OpenStudy (matlee):

168.184.220.11

OpenStudy (anonymous):

2x+6?

OpenStudy (michele_laino):

@Abhisar thank you! :)

OpenStudy (michele_laino):

@CoDsTeR9000 be careful, please you have made a sign error

OpenStudy (anonymous):

-2x-6=-2x-6

OpenStudy (anonymous):

x=All real numbers

OpenStudy (michele_laino):

@CoDsTeR9000 please try to calculate: -2(x+1)=... be careful to sign

OpenStudy (anonymous):

Um -2x-6=-2 (x+1)-4

OpenStudy (michele_laino):

ok! now try to calculate: -2(x+1)=...please

OpenStudy (anonymous):

Sorry if I type wrong im using a tablet

OpenStudy (michele_laino):

don't worry! :)

OpenStudy (anonymous):

-2x6=-2x-(-2)-4

OpenStudy (michele_laino):

I think: -2x-6=-2x-2-4 do you agree?

OpenStudy (anonymous):

Oh missed a negative symbol

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

please note that when you multiply a quantity into a parenthesis, by a negative number, then all number which are into that parenthesis change their sign, namely: -2*(a+b)=-2a-2b

OpenStudy (michele_laino):

so, now perform this calculus: -2-4=...

OpenStudy (michele_laino):

please

OpenStudy (anonymous):

-2-4=2

OpenStudy (michele_laino):



OpenStudy (michele_laino):

plese, retry!

OpenStudy (michele_laino):

oops.. please, retry

OpenStudy (anonymous):

Oooohhhhhh so -2-4=-6 because they're both negative and are... wait.......

OpenStudy (michele_laino):

@CoDsTeR9000